OpenStudy (anonymous):

which is the lowest common denominator, 3/4, 4/7 or 1/2

OpenStudy (anonymous):

the lcd of what?

OpenStudy (saifoo.khan):

Of all these?

OpenStudy (anonymous):

The lowest denominator is 28. because 3*7/7*4 = 21/28 4*4/7*4= 16/28 and 1*14/2*14= 14/28